Your parents gave you $2,000 for your 16th birthday. You want to deposit it in a CD account that is earning 3% annually. Calculate how much your gift will be worth when you graduate from college exactly six years from now using the following methods: a. Long-hand formula (Do not round intermediate calculations. Round final answer to 2 decimal places.) b. Reference table. Use Table A-1. (Round final answer to 2 decimal places.) c. Financial calculator (Round your answer to 2 decimal places.)
Answer #1

1.
Future Value=Present Value*(1+rate)^time=2000*(1+3%)^6=$2,388.10

2.
From the table, Future Value factor of an Annuity for 6 years at 3% is 1.1941
Hence, gift will be worth=1.1941*2000=$2,388.20

3.
N=6
I/Y=3%
PV=2000
PMT=0
CPT FV=$2,388.10